OpenStudy (anonymous):

[Will give medal+fan] A radiator contains 5 quarts of fluid, 20% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 55% antifreeze? (Round your answer to two decimal places.)

OpenStudy (kropot72):

Amount of pure antifreeze in 55% antifreeze fluid = 5 * 0.55 = 2.75 quarts. Let the amount to be drained be x quarts. Therefore the amount of pure antifreeze needed to be added is also x quarts. The amount of 20% antifreeze remaining after draining is 5 - x quarts. Now we can write an equation based on amounts of pure antifreeze as follows: 0.2(5 - x) + x = 2.75 which can be simplified to: x - 0.2x = 2.75 - 1 ..........(1) Now you just need to solve equation (1) to find how much fluid should be drained and replaced with pure antifreeze, which is the value of x.